# The role of magnetic fields in shaping $\gamma$-ray emission from the Fermi bubbles

### v1

@@ 1 Introduction
In 2010, the discovery of large-scale diffuse Galactic emission in the form of two lobes extending above and below the Galactic centre was reported (Su et al., 2010; Dobler et al., 2010), utilising data from the $\gamma-$ray satellite Fermi-LAT. These features indicate the presence of a Galactic-scale outflow, extending up to $\sim$10 kpc above the Galactic plane, which are now known as the Fermi bubbles. Some key features of this emission are a near-constant surface brightness intensity and relatively sharp edges (Ackermann et al., 2014). The width of the base of the Fermi bubbles appears to coincide with the microwave haze (Finkbeiner, 2004; Planck Collaboration et al., 2013). Moreover, recent observations made in the X-ray band have revealed even larger bubbles, called the eROSITA bubbles, which enshroud the Fermi bubble structures (Predehl et al., 2020), and have been suggested to also originate from a Galactocentric outflow (Zhang et al., 2024a).
@@
To date, no consensus regarding the physical mechanism responsible for producing the Fermi bubbles has yet been reached (for a review, see Sarkar, 2024). Several candidate mechanisms, however, have been proposed. Hadronic $\gamma$-ray emission wind models (Crocker & Aharonian, 2011; Mou et al., 2014, 2015; Sarkar et al., 2017) and leptonic jet models (Guo & Mathews, 2012; Yang et al., 2012; Yang et al., 2013; Yang & Ruszkowski, 2017; Guo, 2017), have considered the existence of a supersonic Galactic outflow out at Fermi bubble size scales (though see Mertsch & Sarkar, 2011; Crocker et al., 2015; Shimoda & Asano, 2024, as exceptions to this). However, the velocity profile of the gas for a supersonic outflow scenario appears to be incompatible with the slowly decreasing outflow velocities inferred using UV absorption line techniques.
In addition to the velocity profile, the gas density profile provides information on the nature of the outflow. A supersonic outflow implies a rapid decrease of gas density with increasing distance in the Galactic halo (Sarkar et al., 2017, 2021) due to the dominance of ram pressure over thermal pressure in the outflow. However, recent observations of O VII absorption measurements and ram pressure stripping observations (Martynenko, 2022), Sunyaev-Zeldovich measurements (Bregman et al., 2022), and thermal X-ray emission (Zhang et al., 2024b), have provided new insights on the Galactic halo density distribution. These results all find consistency with expectations for the halo being supported by thermal pressure in hydrostatic equilibrium (Faerman et al., 2017; Prochaska & Zheng, 2019).
A third model, consistent with the density distribution of the Galactic halo gas being in approximate hydrostatic equilibrium, considers a subsonic outflow model, called a Galactic breeze (Taylor & Giacinti, 2017). Building on this, the propagation of a thermally driven outflow reproducing a subsonic solution (Bondi, 1952; Parker, 1958) was simulated using a hydrodynamic model (Tourmente et al., 2023). This work showed that a Galactic breeze solution is capable of producing a Galactic bubble with a decelerating velocity profile, giving rise to $\gamma$-ray emission which broadly matches the Fermi-LAT observations (Ackermann et al., 2014). However, in this model the $\gamma$-ray emission was produced in a broader region than the Fermi bubble $\gamma$-ray emission.
Magnetic fields in the outflow may alter several aspects of a hydrodynamic outflow description. Indeed, observations of Galactic synchrotron emission indicate that magnetic fields permeate the Fermi bubbles, with a strength of 6-12 $\mu$G for the volume-filled lobes and 13-15 $\mu$G for the case of shell-dominated emission edges (Carretti et al., 2013). Recent simulations on the deflection of ultra-high energy cosmic rays predict a magnetic field strength of $\sim$ 7 $\mu$G within a spatially extended region of $\sim$ 6 kpc above the Galactic centre (Shaw et al., 2022). Here, we extend our previous hydrodynamic model to a magnetohydrodynamic (MHD) model and investigate the effects introduced by the presence of magnetic fields.
@@ 2 Galactic outflow model
$\frac{\partial}{\partial t}\left(\rho\textbf{{v}}\right)+\bm{\nabla}\cdot\left(\rho\textbf{{v}}\textbf{{v}}+P\textbf{{I}}\right)-\left(\frac{\textbf{B}\textbf{B}}{4\pi}-\frac{B^{2}\textbf{I}}{8\pi}\right)=-\rho\bm{\nabla}\Phi,$ (2)
$\frac{1}{c}\frac{\partial\textbf{{B}}}{\partial t}+\bm{\nabla}\times\left(\bm{v}\times\bm{B}\right)=0.$ (3)
where $\rho$ is the mass density, v is the velocity vector, $S_{\rho}$ is a source term representing the mass injection rate (see section 2.1.2), $P$ is the thermal pressure, I is the unitary tensor, B is the magnetic field vector, and $c$ is the speed of light. The effect of the total Galactic gravitational potential, $\Phi$ (see section 2.1.1), is represented by the term on the right-hand side of Eq. (2). On the left-hand side of Eq. (2), the second term corresponds to the ram and thermal pressure. The third term corresponds to the magnetic tension and the magnetic pressure. A second-order linear spatial reconstruction with a van Leer limiter has been used in combination with a second-order Runge-Kutta scheme to advance the equations in time. An HLL Riemann solver (Miyoshi & Kusano, 2005) is used for computing the intercell fluxes. The divergence-free condition for the magnetic field is maintained using the divergence cleaning scheme (Dedner et al., 2002). As the simulations have been considered for an isothermal Galactic halo (see Section 2.1), we omit here the energy continuity equation.
@@ 2.1.1 Galactic gravitational potential and density distribution
Figure 2: Spatial distribution of the steady-state (a) velocity and (b) magnetic field strength, both for a 2D map with cylindrical coordinates. The white arrows represent the direction of the respective vector field. The continuous colour bar indicates the logarithm in base 10 for each quantity. For Fig. 2 the black contour lines represent the velocity distribution for two specific values of 100 km s-1 and 30 km s-1. For Fig. 2 the black contour lines represent magnetic field strengths of 0.1 $\mu$G and 0.03 $\mu$G. Fig. 2 shows the distribution of the velocity as a function of Galactic height, $z$, for $R=300$ pc. The solid orange line represents the velocity profile from the MHD simulation, and the dashed blue line represents the velocity profile for the hydrodynamic solution. The dotted red line represents the thermal velocity. Fig. 2 shows the distribution of both magnetic field components, $B_{r}$ and $B_{\phi}$, along the cylindrical radius, $R$, for $z=300$ pc. The solid blue and orange lines represent the distribution of $B_{r}$ and $B_{\phi}$. For comparison, the dashed green line represents the analytic solution for the distribution of $B_{\phi}$.
@@
$n_{\mathrm{gas}}=n_{10}\exp{\left(-\frac{\Phi}{c^{2}_{s}}\right)},$ (4)
where $n_{\rm gas}=\rho/m_{p}$, and $n_{10}=3\times 10^{-3}$ cm-3 is the number density at a distance, $r_{10}=10$ kpc from the Galactic centre. This value has been chosen to match with observations of the O VII spectrum and ram pressure stripping (Martynenko, 2022), as shown in Fig. 1. The thermal velocity, $c_{s}$, is chosen to be 250250250 km s-1, corresponding to an energy of $kT\sim 400$ eV.
@@ 2.1.2 Boundary conditions
The initial radial velocity corresponds to the value calculated analytically to reach a maximum Mach number of M=0.85 at a Galactic radius of 1 kpc. This is $v_{r_{0}}=211$ km s-1, the same as that adopted for our earlier hydrodynamic model (Tourmente et al., 2023). For the initial azimuthal velocity, $\sim$10of the rotational velocity of the Galactic disc has been chosen, corresponding to $v_{\phi_{0}}=20$ km $\rm{s}^{-1}$ (Honma et al., 2012, 2015). The initial value for $\dot{M}$ depends on the initial value of both the velocity and gas density. The chosen value for $v_{r_{0}}$ corresponds to $\dot{M}=0.75~M_{\odot}$ yr-1. This value lies in the upper range of the estimated star formation rate, $0.03-1M_{\odot}$ yr-1, at the Galactic centre (Henshaw et al., 2023). In combination with the initial value for the velocity, the injected kinetic energy luminosity can be calculated, for which one obtains $L_{E_{k}}\approx 10^{40}$ erg s-1.
For our simulations the initial (inner radial) magnetic field values of $B_{r_{0}}$ and $B_{\phi_{0}}$ have been chosen so as to not significantly disrupt the velocity profile. For the radial magnetic field an initial value of $B_{r_{0}}=20~\mu$G has been chosen. The initial value for the azimuthal magnetic field, $B_{\phi_{0}}=2~\mu$G, has been chosen. Further consideration of the role of magnetic fields in Eq. 2 is left to appendix A. In particular, it is highlighted that $B_{\phi}$ can potentially influence the velocity profile. This happens through the magnetic tension term when the magnetic field lines, initially pointing to the $r$-direction, bend over and point in the $\phi$-direction.
@@ 2.2 MHD simulation results
The two panels in Fig. 2 show the velocity and magnetic field distributions in steady-state out to $r=10$ kpc obtained from the numerical outflow simulation with the setup presented in section 2.1. The colour bar in Figs. 2-2 denotes the logarithm of the velocity and magnetic field strengths, respectively. Fig. 2 presents the 2D velocity distribution as a function of the cylindrical coordinates ($R$, $z$). Two contour lines (solid black lines) have been drawn, one for a velocity of 100 km s-1 and one for 30 km s-1. These MHD results produce a velocity profile very similar to our previously considered hydrodynamical model results (Tourmente et al., 2023, section 4.1). The velocity reaches a maximum of 210$\sim$ 210∼ 210 km s-1 at a Galactic radius of 1 kpc and slowly decelerates continuously beyond this point.
Fig. 2 presents the 2D magnetic field distribution as a function of position in cylindrical coordinates ($R$, $z$). The two contour lines shown in the figure (dashed black lines) correspond to 0.1 $\mu$G and 0.03 $\mu$G. For the first 10 kpc from the Galactic centre, the magnetic field strength decreases continuously with increasing distance. Further out, a shell structure is formed at a distance of 12-14 kpc from the Galactic centre, i.e., beyond the range of the Fermi bubbles. The formation of this shell is due to the values chosen for $v_{r_{0}}$, $v_{\phi_{0}}$, $B_{r_{0}}$, and $B_{\phi_{0}}$.
Fig. 2 and Fig. 2 show the velocity and magnetic field, respectively. As the Fermi bubbles extend mainly along the $z$-direction, Fig. 2 emphasises the velocity distribution along the Galactic height, $z$, for $R=300$ pc. Fig. 2 emphasises the $B_{r}$ and $B_{\phi}$ distribution along the cylindrical radius for $z=300$ pc. For Fig. 2, the velocity profile produced by the MHD model is shown by the solid orange line, with the dashed blue line showing the corresponding result obtained for the hydrodynamic model (Tourmente et al., 2023). As appreciated from the similarity of these results, the MHD result describes a subsonic velocity profile broadly similar to that obtained for the hydrodynamical model.
@@
where $\alpha$ is a power-law exponent that depends on the slope of $v_{r}(r)$. The magnetic field is anchored to the outflow that will be influenced in turn by the magnetic tension or magnetic pressure. From Fig. 2 it can be seen that, out to a radius of 10 kpc from the Galactic centre, the distribution of $B_{\phi}$ obtained with the MHD simulation matches well with the analytical distribution (dashed green line) obtained with Eq. (5). Beyond this distance the magnetic field forms a shell, leading to a departure of the MHD magnetic field from the analytic expression at larger distances. However, it should be noted that for our simulations this shell is not in steady-state at the end of the simulation time and is still evolving.
@@ 3 Cosmic ray transport model
$\frac{\partial f}{\partial t}=\mbox{\boldmath$\nabla$}\cdot\left(\mbox{\boldmath$D$}\cdot\nabla f-\mbox{\boldmath$v$}f\right)+\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[\left(\mbox{\boldmath$\nabla$}\cdot\mbox{\boldmath$v$}\right)\frac{p^{3}}{3}f\right]-\frac{f}{\tau_{\rm loss}}+\frac{Q}{p^{2}},$ (6)
where $f$ is the CR space phase density, given by $f=\frac{dN}{d^{3}xd^{3}p}$, where $dN$ is the number of particles within a volume of size $d^{3}xd^{3}p$. The outflow velocity, v, is obtained from the MHD simulation (see section 2), $p$ is the CR momentum, $\tau_{\rm loss}$ is the CR energy loss time scale through inelastic proton-proton ($pp$) collisions (see section 3.3) and $Q$ represents the number of CRs injected per unit of time and per unit of volume (see section 3.1). $D$ is the anisotropic diffusion tensor and is expressed as,
[…]
@@
Since the MHD simulation grid range started from radius $r_{0}$ (300 pc), an interpolation of the velocity and magnetic field distribution was initially made onto the cylindrical CR transport grid. For radii on the CR grid that are smaller than those on the MHD grid (i.e., $<r_{0}$), the velocity and magnetic fields obtained from the MHD grid were extrapolated onto the CR grid.
The CR momentum grid ranges from 10 GeV/c to 30 GeV/c with 5 logarithmically spaced momentum bins. The simulations have been run for a simulation time of 300 Myr, i.e., until a steady-state is reached out to a radial distance of $r\approx 10$ kpc. This distance corresponds approximately to the height reached by the Fermi bubbles. The inner spatial boundaries, $R_{\mathrm{in}}$ and $z_{\mathrm{in}}$, have both been set up with a reflective boundary condition. For the outer spatial boundaries, $R_{\mathrm{out}}$ and $z_{\mathrm{out}}$, both have been set up with an outflow boundary condition.
Simulations of CR transport have been performed for only positive $z$. For Figs. 4.1 and 5, showing CR density distribution and $\gamma$-ray emission for both the north and south Galactic hemispheres in sections 4 (see also appendix D and E) the results have been reflected in the $z=0$ axis since we assume the system is symmetric about the $z$-axis, an equivalent set of results would also be expected for the negative $z$ region.

### v2

@@ 1 Introduction
In 2010, the discovery of large-scale diffuse Galactic emission in the form of two lobes extending above and below the Galactic centre was reported (Su et al., 2010; Dobler et al., 2010), utilising data from the $\gamma$-ray satellite Fermi-LAT. These features indicate the presence of a Galactic-scale outflow, extending up to $\sim$10 kpc above the Galactic plane, which are now known as the Fermi bubbles. Some key features of this emission are a near-constant surface brightness intensity, relatively sharp edges of the emission, and an approximately power-law spectral energy distribution described by $dN/dE\propto E^{-1.9\pm 0.2}$ (Ackermann et al., 2014). The width of the base of the Fermi bubbles appears to coincide with the microwave haze (Finkbeiner, 2004; Planck Collaboration et al., 2013). Moreover, recent observations made in the X-ray band have revealed even larger bubbles, called the eROSITA bubbles, which enshroud the Fermi bubble structures (Predehl et al., 2020), and have been suggested to also originate from a Galactocentric outflow (Zhang and others, 2024a).
@@
To date, no consensus regarding the physical mechanism responsible for producing the Fermi bubbles has yet been reached (for a review, see Sarkar, 2024). Several candidate mechanisms, however, have been proposed. Hadronic $\gamma$-ray emission wind models (Crocker and Aharonian, 2011; Mou et al., 2014, 2015; Sarkar et al., 2017) and leptonic jet models (Guo and Mathews, 2012; Yang et al., 2012, 2013; Yang and Ruszkowski, 2017; Guo, 2017), have primarily focused on the existence of a supersonic Galactic outflow out at Fermi bubble size scales (though see Mertsch and Sarkar, 2011; Crocker et al., 2015; Shimoda and Asano, 2024, as exceptions to this). However, the velocity profile of the gas for a supersonic outflow scenario appears to be incompatible with the slowly decreasing outflow velocities inferred using UV absorption line techniques.
In addition to the velocity profile, the gas density profile provides information on the nature of the outflow. A supersonic outflow implies a rapid decrease of gas density with increasing distance in the Galactic halo (Sarkar et al., 2017, 2021) due to the dominance of ram pressure over thermal pressure in the outflow. However, recent observations of O VII absorption measurements and ram pressure stripping observations (Martynenko, 2022), Sunyaev-Zeldovich measurements (Bregman et al., 2022), and thermal X-ray emission (Zhang and others, 2024b), have provided new insights on the Galactic halo density distribution. These results all find consistency with expectations for the halo being supported by thermal pressure in hydrostatic equilibrium (Faerman et al., 2017; Prochaska and Zheng, 2019).
A third model, consistent with the density distribution of the Galactic halo gas being in approximate hydrostatic equilibrium, considers a subsonic outflow model, called a Galactic breeze (Taylor and Giacinti, 2017). Building on this, the propagation of a thermally driven outflow reproducing a subsonic solution (Bondi, 1952; Parker, 1958) was simulated using a hydrodynamic model (Tourmente et al., 2023). This work showed that a Galactic breeze solution is capable of producing a Galactic bubble with a decelerating velocity profile, giving rise to $\gamma$-ray emission which broadly matches the Fermi-LAT observations (Ackermann et al., 2014). However, in this model the $\gamma$-ray emission was produced in a broader region than the Fermi bubble $\gamma$-ray emission.
[…]
@@ 2 Galactic outflow model
$\frac{\partial}{\partial t}\left(\rho\textbf{{v}}\right)+\bm{\nabla}\cdot\left(\rho\textbf{{v}}\textbf{{v}}+P\textbf{{I}}\right)-\nabla\cdot\left(\frac{\textbf{B}\textbf{B}}{4\pi}-\frac{B^{2}\textbf{I}}{8\pi}\right)=-\rho\bm{\nabla}\Phi,$ (2)
$\frac{\partial\textbf{{B}}}{\partial t}-\bm{\nabla}\times\left(\bm{v}\times\bm{B}\right)=0.$ (3)
where $\rho$ is the mass density, v is the velocity vector, $S_{\rho}$ is a source term representing the mass injection rate (see Section 2.1.2), $P$ is the thermal pressure, I is the unitary tensor, B is the magnetic field vector, and $c$ is the speed of light. The effect of the total Galactic gravitational potential, $\Phi$ (see Section 2.1.1), is represented by the term on the right-hand side of Eq. (2). On the left-hand side of Eq. (2), the second term corresponds to the ram and thermal pressure. The third term corresponds to the magnetic tension and the magnetic pressure. A second-order linear spatial reconstruction with a van Leer limiter has been used in combination with a second-order Runge-Kutta scheme to advance the equations in time. An HLL Riemann solver (Miyoshi and Kusano, 2005) is used for computing the intercell fluxes. The divergence-free condition for the magnetic field is maintained using the divergence cleaning scheme (Dedner et al., 2002). We assume, as a first approximation, that the gas in the Galactic halo has an isothermal temperature distribution (see Section 2.1.1), and we therefore omit the energy continuity equation.
@@ 2.1.1 Galactic gravitational potential and density distribution
$n_{\mathrm{gas}}=n_{10}\exp{\left(-\frac{\Phi-\Phi_{10}}{c^{2}_{s}}\right)},$ (4)
where $n_{\rm gas}=\rho/m_{p}$, and $n_{10}=3\times 10^{-3}$ cm-3 is the number density at a distance, $r_{10}=10$ kpc from the Galactic centre. $\Phi_{10}$ represents the value of the gravitational potential at $r=10$ kpc. This value has been chosen to match with observations of the O VII spectrum and ram pressure stripping (Martynenko, 2022), as shown in Fig. 1. As the Galactic halo is approximated with an isothermal temperature distribution, the thermal velocity, $c_{s}$, is constant and is chosen to be $250$ km s-1, corresponding to an energy of $kT\sim 400$ eV. The limitations of this model assumption are further discussed in Section 5. A density distribution for the Galactic disc has not been included in our initial condition setup, but its gravitational influence is present through $\Phi_{\rm d}$. The scale height for the Galactic disc is 300 pc (Tourmente et al., 2023, Table 1). This is also the position of our inner radial boundary of the simulations, which means that the effect of neglecting the Galactic disc density should be negligible. The effect of neglecting the disc gas density in our initial condition setup is discussed in Section 3.1.
@@ 2.1.2 Boundary conditions
The initial radial velocity corresponds to the value calculated analytically to reach a maximum Mach number of M=0.85 at a Galactic radius of 1 kpc. This is $v_{r_{0}}=211$ km s-1, the same as that adopted for our earlier hydrodynamic model (Tourmente et al., 2023). For the initial azimuthal velocity, $\sim$10% of the rotational velocity of the Galactic disc has been chosen, corresponding to $v_{\phi_{0}}=20$ km $\rm{s}^{-1}$ (Honma et al., 2012, 2015). The initial value for $\dot{M}$ depends on the initial value of both the velocity and gas density. The chosen value for $v_{r_{0}}$ corresponds to $\dot{M}=0.75~M_{\odot}$ yr-1. This value lies in the upper range of the estimated star formation rate, $0.03-1M_{\odot}$ yr-1, at the Galactic centre (Henshaw et al., 2023). In combination with the initial value for the velocity, the injected kinetic energy luminosity can be calculated, for which one obtains $L_{E_{k}}\approx 10^{40}$ erg s-1.
[…]
@@ 2.2 MHD simulation results
The two panels in Fig. 2 show the velocity and magnetic field distributions in steady-state out to $r=10$ kpc, obtained from the numerical outflow simulation with the setup presented in Section 2.1. The colour bar in Figs. 2-2 denotes the logarithm of the velocity and magnetic field strengths, respectively. Fig. 2 presents the 2D velocity distribution as a function of the cylindrical coordinates ($R$, $z$). Two contour lines (solid black lines) have been drawn, one for a velocity of 100 km s-1 and one for 30 km s-1. The velocity reaches a maximum of $\sim 210$ km s-1 at a Galactic radius of 1 kpc and slowly decelerates continuously beyond this point.
[…]
Fig. 2 and Fig. 2 show the velocity and magnetic field, respectively. For both Fig. 2 and Fig. 2, the components $v_{\theta}$ and $B_{\theta}$ are negligible in comparison of $v_{r}$, $v_{\phi}$, $B_{r}$, and $B_{\phi}$, respectively, and have therefore not been included. As the Fermi bubbles extend mainly along the $z$-direction, Fig. 2 emphasises the velocity distribution along the Galactic height, $z$, for $R=300$ pc. Fig. 2 emphasises the $B_{r}$ and $B_{\phi}$ distribution along the cylindrical radius for $z=300$ pc. For Fig. 2, the velocity profile produced by the MHD model is shown by the solid orange line, with the dashed blue line showing the corresponding result obtained for the hydrodynamic model (Tourmente et al., 2023). As appreciated from the similarity of these results, the MHD result describes a subsonic velocity profile broadly similar to that obtained for the hydrodynamical model.
[…]
@@
where $\alpha$ is a power-law exponent that depends on the slope of $v_{r}(r)$. The magnetic field is anchored to the outflow that will be influenced in turn by the magnetic tension or magnetic pressure.
As seen in Fig. 2, out to a radius of 10 kpc from the Galactic centre, the distribution of $B_{\phi}$ obtained with the MHD simulation matches well with the analytical distribution (dashed green line) obtained with Eq. (5). A shell structure is formed at a distance of 12-14 kpc from the Galactic centre, out at the radius where $B_{\phi}$ becomes larger than $B_{r}$. At radii larger than this distance, the magnetic field is azimuthal in direction, and the velocity flow is no longer along the magnetic field lines. This misalignment of the velocity flow and magnetic field results in the primarily the advection, and a lesser extent the compression, of the azimuthal magnetic field, resulting in the formation of the azimuthal magnetic field shell. Appendix C presents the influence of the compression and advection terms of the induction equation that lead to the formation of the shell. However, it should be noted that for our simulations this shell is not in steady-state at the end of the simulation time and is still evolving.
@@ 3 Cosmic ray transport model
$\frac{\partial f}{\partial t}=\mbox{$\nabla$}\cdot\left(\mbox{$D$}\cdot\nabla f-\mbox{$v$}f\right)+\frac{1}{p^{2}}\frac{\partial}{\partial p}\left[\left(\mbox{$\nabla$}\cdot\mbox{$v$}\right)\frac{p^{3}}{3}f\right]-\frac{f}{\tau_{\rm loss}}+\frac{Q}{p^{2}},$ (6)
@@
Since the MHD simulation grid range started from radius $r_{0}$ (300 pc), an interpolation of the velocity and magnetic field distribution was initially made onto the cylindrical CR transport grid. For radii on the CR grid that are smaller than those on the MHD grid (i.e., $r_{0}$), the velocity and magnetic fields obtained from the MHD grid were extrapolated onto the CR grid.
The CR momentum grid ranges from 10 GeV/c to 30 GeV/c with 5 logarithmically spaced momentum bins. The simulations have been run for a simulation time of 300 Myr, i.e., until a steady-state is reached out to a radial distance of $r\approx 10$ kpc. This distance corresponds approximately to the height reached by the Fermi bubbles. The inner spatial boundaries, $R_{\mathrm{in}}$ and $z_{\mathrm{in}}$, have both been set up with a reflective boundary condition. For the outer spatial boundaries, $R_{\mathrm{out}}$ and $z_{\mathrm{out}}$, both have been set up with an outflow boundary condition. The inner and outer momentum boundaries have been set up as outflow (Rodgers-Lee et al., 2020).
Simulations of CR transport have been performed for only positive $z$. For Figs. 4.1 and 5, showing CR density distribution and $\gamma$-ray emission for both the north and south Galactic hemispheres in Section 4 (see also appendix E and F) the results have been reflected in the $z=0$ axis since we assume the system is symmetric about the $z$-axis, an equivalent set of results would also be expected for the negative $z$ region.